3.   You are HR director for a growing law firm in Boise, Idaho, which currently has need of writing 48 legal briefs every hour. Each of your company’s attorneys can write on average eight briefs per hour. You are considering hiring four paralegals to shoulder the load; each paralegal is slower than the attorneys and can write on average only four briefs per hour. You scan the current wages in the Boise area and notice that the lawyers in your company earn the local occupational median wage of $44.05 per hour, but that the prospective four paralegals will likely want to get paid their local occupational median wage of $24.86 per hour.

a.   Would your company save money in the writing of the 48 legal briefs by hiring the four new paralegals and firing some attorneys?

b. The Lawyers of Idaho Association of Research Scientists has observed that the Bureau of Labor Statistics projects that employment of paralegals over the next decade will rise almost twice as fast as the employment of lawyers, which it assumes will push paralegal wages down to $21.25 due to their (relatively higher) labor supply. If the other values remain the same (attorney wage and both writing speeds, need for 48 briefs per hour), would the company save money in the writing of the 48 legal briefs by hiring four paralegals and firing some attorneys?
Answer #1

a.)

Four paralegals can write 16 legal briefs every hour.

No of attorneys. needed to write 16 legal briefs every hour = 16/8 = 2 attorneys.

So if we hire 4 paralegals, we can at most fire 2 attorneys in place.

If we hire 4 paralegal, we pay = $24.86 * 4 = $99.44 per hour

If we fire 2 attorneys, we gain = $44.05*2 = $88.1 per hour

Thus net gain by hiring 4 paralegals= $88.1 -$99.44 =  -\$ 11.34 per hour

Thus, the company would not save any money by hiring 4 paralegals and firing some attoneys. Instead it will lose $11.34.

b)

If the wages of paralegal is reduced to $22.15 per hour,

The total wages of 4 paralegal = $22.15*4 = $88.60 per hour

So, if we fire 2 attorneys, we gain $88.1 per hour.

Thus net gain by hiring 4 paralegals = $88.60 - $88.1 = $0.50 per hour.

Yes, the company would save money($.50 per hour) in the writing of the 48 legal briefs by hiring four paralegals and firing some attorneys.